OpenStudy (anonymous):

how you you solve 5√320? using perfect and non perfect squares?

OpenStudy (anonymous):

5sqrt(320)=5sqrt(4*8*10) =5sqrt(4*4*4*5) =5*8sqrt(5) 40sqrt(5)